In a study of reaction rates, you need to dilute copper(II) sulfate solution. You take 5.00 mL of 0.050 M CuSO4(aq) and dilute this to a final volume of 100.0 mL

a) What is the final concentration of the dilute solution?


b) What mass of CuSO4(s) is present in 10.0 mL of the final dilute solution?

c) Can this final dilute solution of 10 mL be prepared directly using the pure solid? Defend your answer.
1
Expert's answer
2020-06-08T15:41:04-0400

A). (5.00ml)(0.005mol/L) =

C2 = (100.0ML)


C2 =0.000250MOL/L


B). N = C2 = (0.00025)(0.010L)

= (2.5)(10-6 )


M = NM = [ (2.5)(10-6 mol)] (159.62g/mol)


= (3.99)(10- 4 g)


= 0.000399g